# Breaking the Trend: How to Avoid Cherry-Picked Signals

### v3

@@ 4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. 3), introduced in Sec 3 and, therefore, the diffusive process in Eq. 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017). This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers. Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. 1 follows a particular random matrix structure, which is considered realistic. Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
@@ 4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2 which is an very easy solution to reduce trading cost (usually, trend indicators include rules to reduce trading costs. While the $EMA$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective. Benichou at al (2017) recommended optimization under constraints to better manage trading costs).
Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$.
@@ 4.4 Dataset and different simulated parameters
The simulation starts on the 25th of May 1990 and stops on the 7th of December 2023. We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures. The description is in the appendix A.
@@ 5.1 Grebenkov’s model empirical validation
Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$. The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends. The risk of overfitting is very limited, as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index. Using an extended dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value. Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig. 2) applied to the Dow Jones Index, as our fitting is performed on a broader set of asset classes over a more recent period. That is the main result of the paper.
@@
As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab. 4 displays holding periods slower than 80 days). Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum. The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign. Additionally, they did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix—a method that explains a significant part of the Sharpe ratio, as shown in Benichou at al (2017). For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
@@
Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ as signal v.s. $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3. The theoretical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ . Empirical measurements are displayed in the tab 4.

### v4

@@ 4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. 3), introduced in Sec 3 and, therefore, the diffusive process in Eq. 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017). We decided to use the $ARP$ as the projection portfolio, as it is theoretically and empirically considered to be the optimal trend-following portfolio, capable of minimizing measurement noise in the Sharpe ratio which is empirically optimal (Benichou at al (2017)). However, we expect that the results should also hold even with lower Sharpe ratios and more noises for more conventional and naive trend-following portfolios, such as those that do not manage correlations or assume an identity correlation matrix except when computing volatility used to target the size of the portfolio.
ARPARP constructs a rotationally invariant portfolio which makes the Sharpe ratio less dependent on the subjective selection of futures used to define the investment universe, as well as on discretionary choices regarding the relative number of FX, stock index, bond, and commodity futures. $ARP$ also by consequence applies trend-following to eigenvenvectors while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers. Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. 1 follows a particular random matrix structure, which is considered realistic. Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
@@ 4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2 which is an very easy solution to reduce trading cost (usually, trend indicators include rules to reduce trading costs. While the $EMA$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective. Benichou at al (2017) recommended optimization under constraints to limit trading and illiquidity and to better manage trading costs).
Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$. The target-volatility step is both usual and useful to get standardized and homoscedastic returns for the portfolio so that Sharpe ratio is adapted.
We applied exactly the same computation for the naïve approach, except that we assumed $C^{-0.5}$ in Eq. 12 to be the identity matrix. This is equivalent to managing trend-following strategies independently, underlying by underlying, while constraining the overall portfolio to maintain constant volatility. This approach is referred to as ‘naïve’ by Benichou at al (2017). We did not test the true Markowitz portfolio, which would simply replace $C^{-0.5}$ by $C^{-1}$, as that solution performed significantly worse.
@@ 4.4 Dataset and different simulated parameters
The simulation starts on the 29th of May 1990 and stops on the 7th of December 2023. We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures. The description is in the appendix A.
@@ 5.1 Grebenkov’s model empirical validation
Interestingly, the empirical simulation for ARP fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$. The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends. The risk of overfitting is very limited, and the fit was almost out-of sample as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index. Using a 70 times larger dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value. Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig. 2) applied to the Dow Jones Index, as our fitting is performed on a broader set of asset classes over a more recent period. That is the main result of the paper.
We can note that the measurements using $ARP$ fit the theoretical results better than those obtained with the naïve portfolio, where the correlation matrix is not considered to optimize diversification and reduce noise. The naïve measurements are not only shifted downward as we could have expected (the shift could have been more brutal if the universe was not well diversified to avoid potential criticism), but also appear slightly rotated (with $\lambda$ slightly increased) — a difference that may simply result from noisier estimates, or from the assumption made in Section 3 regarding the extension of Grebenkov’s formula, using the same $\lambda$, from the one-dimensional to the multi-dimensional case. It may be that Grebenkov’s formula holds more rigorously, from a mathematical standpoint, for the optimal $ARP$ portfolio than for the naïve one. However, this would be quite surprising, since Grebenkov’s formula is derived precisely under the mathematical conditions that make the naïve approach optimal (correlations between $\xi_{i,k}$ and those between $\epsilon_{i,t}$ are set to zero), as detailed in Section 3. The most likely explanation, which is not very intuitive since the error in the graph is highly autocorrelated with an estimated std of error of 0.18 in line with the 30 years of history, is that the $ARP$ measurements effectively apply a weighted average of $\lambda$ per market, assigning higher weights to markets which are different from other and which are not simply repetitions of others (as in the stock indices cases where all indices are very close). In contrast, the naïve approach uses an equal-weight scheme, which can overweight markets like stock indices that repeat themselves, thereby amplifying measurement errors. As a result, the naïve portfolio has just overweighted markets which appear faster. Schmidhuber (2021) similarly aggregates markets to obtain more robust results, observing that these coefficients are universal across asset classes and exhibit a universal scaling behavior, as the trend’s time horizon ranges from a few days to several years.
@@
As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab. 4 displays holding periods slower than 80 days). Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum. The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign. Additionally, they use the ”naive” approach and did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix which slightly improves the fit, as observed in Fig. 5. For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
@@
Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) or the naive approach applied to $\mathrm{EMA}$ as signal v.s. $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3. The theoretical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ . Empirical measurements for the ARP are displayed in the tab 4.
5.2 Scaling factor empirical validation
We can see that the empirical measurements in Fig. 5 appear less noisy than those reported by Ferreira et al. (2018) for the Dow Jones, where a complex ARMA process was required instead of a simple AR(1). It is clear that aggregating data by projecting onto our optimal trend-following portfolio, rather than analyzing each market individually, is highly effective in reducing noise that could otherwise create misleading evidence of mean reversion at longer time scales, as observed in Moskowitz et al (2012).
We applied a simple empirical protocol to illustrate why analyzing the Sharpe ratio as a function of model parameters at the portfolio level is far more efficient than doing so market by market, as in Ferreira et al. (2018) (see their Figure 6), which was affected by considerable noise. Another advantage of this protocol is that it provides an empirical test of the scaling factor suggested in Section 3, where the Sharpe ratio depends on the size of the universe as $\sqrt{\frac{N}{1+\left(N-1\right)\rho^{2}}}$.
We measured the empirical Sharpe ratios for the $ARP$ using an $EMA$ parameter value of $\frac{1}{120}$—close to the theoretical optimum—across several random universes containing 1, 3, 6, 9, 15, 20, and 27 underlying assets, each randomly selected from the original universe of 70 assets. We conducted 20 random trials for each case.
Fig. 6 shows another impressive fit, although the fitted value of $\bm{\rho}^{2}$ (0.024) appears slightly lower compared with the empirical average of squared correlations (0.056) based on the weekly returns. We can expect through that fit a Sharpe ratio of 1.28 for the whole universe of size 70, a Sharpe ratio of 1.40 for a 140 size and a Sharpe ratio of 1.60 for an infinite universe.
The same figure also shows that the Sharpe ratio increases with the size of the universe, while its standard deviation, theoretically expected to remain stable around 0.2, declines only slightly—likely due to the larger correlations between randomly generated universes when they are larger. We can see that the mean Sharpe ratio is only about 0.2 when $N=1$, the one dimensional case, which is of the same order of magnitude as the noise level. This highlights how diversification and data aggregation—through projection onto the ARPARP—help reduce the noise-to-signal ratio of the measurements.
Figure 6: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ with $\eta=1/120$ as signal v.s. NN the size of the universe and the theoretical scaling factor inspired from Eq. (40) from Grebenkov and Serror (2015). The fitted $\bm{\rho}^{2}$ at 0.024 is just lower than the empirical average of squared correlations.